What are the corresponding degrees of freedom and critical value for a one-way ANOVA with 3 conditions, 5 subjects in each condition, and α = 0.01?

BG: ______________________
WG: ______________________
Total: ______________________
critical value: ______________________
Answer #1

Given k=# of Conditionis = 3 nr 5X 3 = 15 d. f-(BG)= def. (WG) = n k = 12 d.f. (Total = n-1-14 x = 0.01 Critical value = 2 Fa